# OSV: One Step is Enough for High-Quality Image to Video Generation

### v1

@@ Introduction
High Order Sampler Based On Time Travel. As shown in Figure 2, we observed the negative impact of using CFG on the distilled model. Even with smaller CFG weights, the improvement in video generation was minimal. We decided to remove CFG and design a higher-order sampler named Time Travel Sampler (TTS). Suppose the number of sampler steps is set to $k$, corresponding to the time function $t^{k}$, and the number of sampler steps is set to $k+1$, corresponding to the time function $t^{k+1}$. $t^{k+1}$ has one more time step compared to $t^{k}$. Let $t^{k}_{0}=0$ and $t^{k+1}_{0}=0$, we have $t^{k+1}_{i}<t^{k}_{i},t^{k+1}_{i}>t^{k}_{i-1},\forall i\in[1,k]$. During sampling, we set the number of sampler steps to $k$. Observing that the image generation quality improves as the time step $t$ decreases, at step $i$, given the sample $x_{t^{k}_{i}}$, we can first solve for the sampling result at the lower time step $t^{k+1}_{i}$ and then revert to $t^{k}_{i}$ to solve the consistency function again:
@@
$&f_{\theta}(\mathbf{x}_{t^{k}_{i}},t^{k+1}_{i})=c_{\text{skip}}(t^{k+1}_{i})\mathbf{x}_{t_{i}^{k}}+c_{\text{out}}(t^{k+1}_{i})F_{\theta}(\mathbf{x}_{t^{k}_{i}},t^{k+1}_{i}),\\ &\hat{\mathbf{x}}_{t^{k}_{i}}=c_{\text{skip}}(t^{k}_{i})f_{\theta}(\mathbf{x}_{t^{k}_{i}},t^{k+1}_{i})+c_{\text{out}}(t^{k}_{i})F_{\theta}(f_{\theta}(\mathbf{x}_{t^{k}_{i}},t^{k+1}_{i}),t^{k}_{i}).\end{aligned}$
Training Data Resolution. We found that setting different training data sizes for different training stages can accelerate the training process. In the first stage, the quality of the generated videos is evidently related to the dataset size, so we use a resolution of 1024$\times$576. In the second stage, we train the model using a resolution of 576$\times$320.
Experiment
Implementation details. We apply stable video diffusion as the base model for most experiments. we uniformly sample 100 timesteps for training. For better training and evaluation of our method, we utilize OpenWebVid-1M (Nan et al. 2024) for training and validation. We randomly select 1 million videos from OpenWebVid-1M as the training set and 1,000 videos as the test set. In the first stage, we fixed the resolution of the training videos at 1024$\times$768, the FPS at 7, the batch size at 1, and the learning rate at 5e-6. The training was conducted over 2K iterations on 8 NVIDIA H800 GPUs. In the second stage, we fixed the resolution of the training videos at 576$\times$320, the FPS at 7, the batch size at 1, and the learning rate at 5e-6. The training was conducted over 20K iterations on 2 NVIDIA H800 GPUs. All stages use the Adam optimizer (Kingma and Ba 2014). We use Fréchet Gesture Distance (FVD) (Unterthiner et al. 2018) to evaluate our model.
@@
Figure 6: Qualitative generation results. One-step result of OSV with TTS achieves superior video clarity compared to other baselines.
@@
Table 1: Image-to-video performance comparison on the validation set of OpenWebVid-1M. $\dagger$ means our implementations of SF-V (no public weight).
@@
Table 2: Ablation study with OSV. Unless otherwise specified, we set the sampling steps to 1 and use the Time Travel Sampler(TTS).
@@
      576$\times$320       171.15
      1024$\times$576       181.95
(h) Effect of Data Resolution in Stage Two.
Quantitative Experiments
@@
Table 6 illustrates the quantitative comparison of OSV with the strong baseline methods Animate-LCM and SF-V. Observing the negative impact of CFG on the distilled models, we remove CFG from both the Animate-LCM and OSV models. This reduction led to decreased inference time for the student network. OSV significantly outperforms the baseline methods, especially at low steps. Our method is compared with SF-V and AnimateLCM, using different numbers of sampling steps for each method. Additionally, we found that using the high-order sampler in a single step resulted in an FVD of 171.15, compared to an FVD of 181.95 when performing direct inference with two steps. This demonstrates better results than direct two-step inference.
@@
Figure 7 shows the results of our user study, where our model achieved higher clarity and smoothness compared to the teacher.
Qualitative Results
@@
Figure 6 shows the generated results of our method in image-to-video generation, all of which achieved satisfactory performance. The generated results demonstrate that our method effectively adheres to the consistency properties across different inference steps, maintaining similar styles and motions. Other methods either suffer from overexposure issues or blurring due to motion. We exhibit good visual quality and smooth motion with only one step.
@@
Figure 7: User study comparing our distilled model with its teacher and competing distillation baselines. Animate-LCM uses 8 sampling steps, SF-V uses 1 sampling step, OSV uses 1 sampling step with TTS, Teacher uses 25 sampling steps. Supplementary materials provide more generated videos.
Ablation Studies
@@
Effect of Multi-Step Solving. We set the OSV model with 5 solver steps as Baseline-1. To verify the effectiveness of the multi-step solving method, we remove the multi-step solving component and train the OSV model with 1 solver step under the same training settings. As shown in Table 2a, the multi-step solving method achieves higher solving accuracy. Considering the training time for all models is the same, our method also performs better, as illustrated in Figure 5.
@@
Effect of First Stage of Training and Second Stage of Traning. We set the OSV model trained with both Stage 1 and Stage 2 as Baseline-2. To verify the effectiveness of Stage 1 training, we removed Stage 1 training and trained the OSV model with only Stage 2 under the same training settings. As shown in Table 2b, it is evident that performing both Stage 1 and Stage 2 training contributes more to the model’s convergence. Stage 1 training primarily ensures that the student model can still generate detailed content at low steps, which aids in the consistency training of Stage 2 and prevents blurring issues similar to those in Animate-LCM. To verify the effectiveness of the second stage of training, we remove the second stage and train the OSV model with only the first stage under the same training settings. As shown in Table 2c, it is evident that the second stage of training contributes more to refining the generated videos. The consistency training in the second stage further enhances the details of the generated videos. We also find that our TTS sampler is effective for distilling the student model using GAN methods.
@@
Effect of Adversarial Training. We set the OSV model with adversarial loss as Baseline-3. To verify the effectiveness of adversarial distillation, we remove the adversarial loss and train the OSV model with only Huber loss and consistency loss under the same training settings. As shown in Table 2d, adversarial training results in higher generation quality. Using only consistency loss leads to a fitting error between the student model and the teacher model.
Effect of VAE Decoder. We set the OSV model trained with only the first stage as Baseline-4. We add the VAE Decoder from the ADD method, as shown in Figure 3, to the discriminator and trained the OSV model with only the first stage. As shown in Table 2e, we find that adding the VAE Decoder resulted in even worse performance. This indicates that performing adversarial training in the latent space is more beneficial for the discriminator. After removing the Vae Decoder, the training speed on the NVIDIA H800 improved from an average of 4.29 seconds/iteration to 2.61 seconds/iteration. In the supplementary materials, we visualize and analyze the differences between latent space videos and pixel space videos.
@@
Effect of CFG. As shown in Table 2f, we investigate the impact of CFG on video generation by the model. Figure 2 illustrates the overexposure issue caused by adding CFG. It is evident that even reducing the CFG scale still negatively affects the model. Removing CFG not only saves time but also improves the quality of the generated videos.
@@
Effect of Data Resolution in Stage One and Stage Two. As shown in Tables 2g and  2h, we find that different training data Resolution can be used in the first and second stages. For the first stage, we did not introduce the consistency distillation loss, so the quality of the generated videos is related to the dataset size. If we use a dataset size of 576×320, the videos are downsampled quite small, resulting in significant information loss. In the second stage, we enforce the consistency of the student model’s trajectories at different time steps, so the quality of the generated videos depends on the quality of the teacher model’s videos, and a larger video resolution is not required. Additionally, since the computational load in the first stage is much lower than in the second, we use a dataset size of 1024$\times$576 in the first stage and 576$\times$320 in the second. This significantly reduces the model’s distillation time.
Conclusion
@@
Limitations. We observe that the distilled model exhibits less motion and heavily relies on the input images. This is a common phenomenon in distilled models with few-step sampling.

### v2

@@ 4 Method
High Order Sampler Based On Time Travel. As shown in Figure 2, we observe the negative impact of using CFG on the distilled model. Even with smaller CFG weights, the improvement in video generation is minimal. We decide to remove CFG and design a higher-order sampler named Time Travel Sampler (TTS). Suppose the number of sampler steps is set to $k$, corresponding to the time function $t^{k}$, and the number of sampler steps is set to $k+1$, corresponding to the time function $t^{k+1}$. $t^{k+1}$ has one more time step compared to $t^{k}$. Let $t^{k}_{0}=0$ and $t^{k+1}_{0}=0$, we have $t^{k+1}_{i+1}<t^{k}_{i+1},t^{k+1}_{i+1}>t^{k}_{i},\forall i\in[1,k-1]$. During sampling, we set the number of sampler steps to $k$. Observing that the image generation quality improves as the time step $t$ decreases, given the sample $x_{t^{k}_{i+1}}$, we can first solve for the sampling result at the lower time step $t^{k+1}_{i+1}$ and then revert to $t^{k}_{i}$ to solve the consistency function again:
@@
$&f_{\theta}^{t^{k}_{i+1}}=c_{\text{skip}}(t^{k}_{i+1})\mathbf{x}_{t_{i+1}^{k}}+c_{\text{out}}(t^{k}_{i+1})F_{\theta}(\mathbf{x}_{t^{k}_{i+1}},t^{k}_{i+1}),\mathbf{x}_{t^{k+1}_{i+1}}=f_{\theta}^{t^{k}_{i+1}}+\sigma_{t^{k+1}_{i+1}}\hat{\epsilon}_{i+1}^{k+1}\\ &\hat{\epsilon}_{i+1}^{k+1}=(\epsilon_{0}+\frac{(\mathbf{x}_{t^{k}_{i+1}}-f_{\theta}^{t^{k}_{i+1}})}{\sigma_{t^{k}_{i+1}}})/2,\\ &f_{\theta}^{t^{k+1}_{i+1}}=c_{\text{skip}}(t^{k+1}_{i+1})\mathbf{x}_{t_{i+1}^{k+1}}+c_{\text{out}}(t^{k+1}_{i+1})F_{\theta}(\mathbf{x}_{t^{k+1}_{i+1}},t^{k+1}_{i+1}),\\ &\hat{\mathbf{x}}_{t^{k}_{i}}=c_{\text{skip}}(t^{k}_{i+1})(f_{\theta}^{t^{k+1}_{i+1}}+\sigma_{t^{k}_{i+1}}\hat{\epsilon}_{i+1}^{k})+c_{\text{out}}(t^{k}_{i+1})F_{\theta}(\mathbf{x}_{t^{k+1}_{i+1}},t^{k+1}_{i+1})+\sigma_{t_{i}^{k}}\epsilon_{0},\hat{\epsilon}_{i+1}^{k}=(\epsilon_{0}+\frac{(\mathbf{x}_{t^{k+1}_{i+1}}-f_{\theta}^{t^{k+1}_{i+1}})}{\sigma_{t^{k+1}_{i+1}}})/2,\\ \end{aligned}$
where $\epsilon_{0}$ is the random Gaussian noise. Using TTS leads to an increase in NFE, but removing CFG reduces NFE.
@@
Figure 6: Qualitative generation results. One-step results of OSV with TTS achieves superior video clarity compared to other baselines. Please refer to the supplementary material for a better comparison of the videos generated by different models.
5 Experiments
Implementation Details. We apply stable video diffusion as the base model for most experiments. we uniformly sample 100 timesteps for training. For better training and evaluation of our method, we utilize OpenVid-1M [18] for training and validation. We randomly select 1 million videos from OpenVid-1M as the training set and 1,000 videos as the test set. In the first stage, we fix the resolution of the training videos at 1024$\times$768, the FPS at 7, the batch size at 1, and the learning rate at 5e-6. The training is conducted over 2K iterations on 8 NVIDIA H800 GPUs. In the second stage, we fix the resolution of the training videos at 576$\times$320 (in the final 10K iterations, the resolution is increased to 1024$\times$768), the FPS at 7, the batch size at 1, and the learning rate at 5e-6. The training is conducted over 20K iterations on 2 NVIDIA H800 GPUs. We find that higher resolution in training datasets generally yields better results, but the impact varies across different stages, as discussed in Sec. 5.3. All stages use the Adam optimizer [12]. We use FVD [31] to evaluate our model. All models, including OSV, use SVD as the basic model architecture (Unet architecture), initialized with the same pretrained model, sharing no other differences with SVD except when explicitly mentioned. Apart from specifying TTS usage, we uniformly evaluate using the solver from Consistency Model.
@@ 5 Experiments
Table 1: Image-to-video performance comparison on the validation set of OpenVid-1M. $\dagger$ means our implementations of SF-V (no public weight).
@@
Table 2: Ablation study with OSV. Unless stated otherwise, we set the sampling steps to 1 and use the Time Travel Sampler (TTS).
@@
      576$\times$320 (10K) and 1024$\times$576 (10K)       171.15
      1024$\times$576       173.14
5.1 Quantitative Experiments
@@
Table 1 illustrates the quantitative comparison of OSV with the strong baseline methods Animate-LCM and SF-V. Observing the negative impact of CFG on the distilled models, we remove CFG from both the Animate-LCM and OSV models. This reduction led to decreased inference time for the student network. OSV significantly outperforms the baseline methods, especially at low steps. Our method is compared with SF-V and AnimateLCM, using different numbers of sampling steps for each method. Additionally, we find that using the high-order sampler in a single step resulted in an FVD of 171.15, compared to an FVD of 181.95 when performing direct inference with two steps. This demonstrates better results than direct two-step inference.
@@
Figure 7 shows the results of our user study, where our model achieve higher clarity and smoothness compared to the teacher.
5.2 Qualitative Results
@@
Figure 6 shows the generated results of our method in image-to-video generation, all of which achieve satisfactory performance. The generated results demonstrate that our method effectively adheres to the consistency properties across different inference steps, maintaining similar styles and motions. Other methods either suffer from overexposure issues or blurring due to motion. We exhibit good visual quality and smooth motion with only one step.
@@
Figure 7: User study comparing our distilled model with its teacher and competing distillation baselines. For each model, we generate 30 videos across diverse scenarios, ask users to vote for the best-performing model. AnimateLCM with 8 sampling steps, SF-V with 1 sampling step, OSV with 1 sampling step and TTS, and the Teacher model with 25 sampling steps.
5.3 Ablation Studies
@@
Effect of Multi-Step Solving Method. We set the OSV model with 5 solver steps as Baseline-1. To verify the effectiveness of the multi-step solving method, we remove the multi-step solving component and train the OSV model with 1 solver step under the same training settings. As shown in Table 6(a), the multi-step solving method achieves higher solving accuracy. Considering the training time for all models is the same, our method also performs better, as illustrated in Figure 5.
@@
Effect of First Stage of Training and Second Stage of Traning. We set the OSV model trained with both Stage 1 and Stage 2 as Baseline-2. To verify the effectiveness of Stage 1 training, we remove Stage 1 training and train the OSV model with only Stage 2 under the same training settings. As shown in Table 6(b), it is evident that performing both Stage 1 and Stage 2 training contributes more to the model’s convergence. Stage 1 training primarily ensures that the student model can still generate detailed content at low steps, which aids in the consistency training of Stage 2 and prevents blurring issues similar to those in Animate-LCM. To verify the effectiveness of the second stage of training, we remove the second stage and train the OSV model with only the first stage under the same training settings. As shown in Table 6(c), it is evident that the second stage of training contributes more to refining the generated videos. The consistency training in the second stage further enhances the details of the generated videos. We also find that our TTS sampler is effective for distilling the student model using GAN methods.
@@
Effect of Adversarial Training. We set the OSV model with adversarial loss as Baseline-3. To verify the effectiveness of adversarial distillation, we remove the adversarial loss and train the OSV model with only Huber loss and consistency loss under the same training settings. As shown in Table 6(d), adversarial training results in higher generation quality. Using only consistency loss leads to a fitting error between the student model and the teacher model.
Effect of VAE Decoder. We use Baseline-3. We add the VAE Decoder from the ADD method, as shown in Figure 3. As shown in Table 6(e), we find that adding the VAE Decoder resulted in even worse performance. This indicates that performing adversarial training in the latent space is more beneficial for the discriminator. As shown in Figure 8, we visualize the latent space and pixel space of the input images. It can be observed that when the input images are compressed by the VAE, the outlines are preserved, retaining a significant amount of low-frequency information. This is beneficial for ViT-like models in feature extraction. We upsample the latent space data to a size suitable for DINOv2 feature extraction using sub-pixel convolution.
@@
Effect of CFG. As shown in Table 6(f), we investigate the impact of CFG on video generation by the model. Figure 2 illustrates the overexposure issue caused by adding CFG. It is evident that even reducing the CFG scale still negatively affects the model. Removing CFG not only saves time but also improves the quality of the generated videos.
@@
Effect of Data Resolution in Stage One and Stage Two. In the first stage, we do not introduce the consistency distillation loss, so the quality of the generated videos relates to the dataset size. If we use a dataset size of 576$\times$320, the videos are downsampled quite small, resulting in significant information loss. In the second stage, we enforce the consistency of the student model’s trajectories at different time steps, so the quality of the generated videos depends on the quality of the original videos and the videos generated by the teacher model. Using a dataset size of 576$\times$320 in the early stages significantly reduces the model’s distillation time and shows little difference in the FVD metric during the second stage. However, we observe that although the FVD metric shows little difference, the SVD model distilled on the low-resolution dataset encounters more failure cases during inference, such as generating videos with smaller motions. Therefore, we recommend training on a high-resolution dataset if sufficient computational resources are available.
Figure 8: Visualize the latent space and pixel space of the input images. To map the latent space to the RGB color space, we quantize the data within the latent space to the range [0, 255].
6 Conclusion
@@
Limitations. There are some bad cases when the distilled model generates human motion. For example, the distilled model produce significant blurring when attempting to generate hand movements. Increasing the number of inference steps to 4 resolve the issue. In future work, we plan to introduce stronger feature extraction networks as replacements for DINOv2. We observe that the distilled model exhibits less motion and heavily relies on the input images. This is a common phenomenon in distilled models with few-step sampling.
Appendix
[…]
Appendix A Proofs
The following is based on consistency distillation [30].
A.1 Multi-Step Solving Method
Theorem A.1.
Let $\Delta t\coloneqq\max_{n\in\llbracket 1,N-1\rrbracket}\{|t_{n+1}-t_{n}|\}$, and $f(\cdot,\cdot;\phi)$ be the target phased consistency function induced by the pre-trained diffusion model (empirical PF-ODE). Assume $f_{\theta}$ satisfies the Lipschitz condition: there exists $L>0$ such that for all $t\in[\epsilon,T]$, $x$, and $y$, we have $\lVert f_{\theta}(x,t)-f_{\theta}(y,t)\rVert_{2}\leq L\lVert x-y\rVert_{2}$. Assume further that for all $n\in\llbracket 1,N-1\rrbracket$, the ODE solver called at $t_{n+1}$ has local error uniformly bounded by $O((t_{n+1}-t_{n})^{p+1})$ with $p\geq 1$. Then, if $Dis(f_{\theta}(x_{t_{n+m}},t_{n+m}),f_{\theta}(\hat{x}_{t_{n}}^{\phi},t_{n}))=0$, we have
$\sup_{n,x}\|f_{\theta}(x,t_{n})-f(x,t_{n};\phi)\|_{2}=O((\Delta t)^{p}).$
Proof.
From the loss $Dis(f_{\theta}(x_{t_{n+m}},t_{n+m}),f_{\theta}(\hat{x}_{t_{n}}^{\phi},t_{n}))=0$, we have:
$f_{\theta}(x_{t_{n+m}},t_{n+m})\equiv f_{\theta}(\hat{x}_{t_{n}}^{\phi},t_{n}).$ (14)
Let $e_{n}\coloneqq f_{\theta}(x_{t_{n}},t_{n})-f(x_{t_{n}},t_{n};\phi).$ We obtain the subsequent recursive formula:
$e_{n+m}$ $=f_{\theta}(x_{t_{n+m}},t_{n+m})-f(x_{t_{n+m}},t_{n+m};\phi)$
=$\stackrel{{\scriptstyle(i)}}{{=}}f_{\theta}(\hat{x}_{t_{n}}^{\phi},t_{n})-f(x_{t_{n}},t_{n};\phi)$
$=f_{\theta}(\hat{x}_{t_{n}}^{\phi},t_{n})-f_{\theta}(x_{t_{n}},t_{n})+f_{\theta}(x_{t_{n}},t_{n})-f(x_{t_{n}},t_{n};\phi)$
$=f_{\theta}(\hat{x}_{t_{n}}^{\phi},t_{n})-f_{\theta}(x_{t_{n}},t_{n})+e_{n},$ (15)
where (i) is due to Eq. 14 and $f(x_{t_{n+m}},t_{n+m};\phi)=f(x_{t_{n}},t_{n};\phi)$. Considering $f_{\theta}(\cdot,t_{n})$ has Lipschitz constant $L$, we have:
$\lVert e_{n+m}\rVert_{2}$ $\leq\lVert e_{n}\rVert_{2}+L\lVert\hat{x}_{t_{n}}^{\phi}-x_{t_{n}}\rVert_{2}$ (16)
=$\stackrel{{\scriptstyle(i)}}{{=}}\lVert e_{n}\rVert_{2}+L\cdot O(\max_{k\in\llbracket n,n+m-1\rrbracket}(t_{k+1}-t_{k})^{p+1})$ (17)
$=\lVert e_{n}\rVert_{2}+O(\max_{k\in\llbracket n,n+m-1\rrbracket}(t_{k+1}-t_{k})^{p+1}).$ (18)
Considering the definition of $f$, we have:
$e_{0}$ $=f_{\theta}(x_{t_{0}},t_{0})-f(x_{t_{0}},t_{0};\phi)$ (19)
=$\stackrel{{\scriptstyle(ii)}}{{=}}x_{t_{0}}-x_{t_{0}}$ (20)
$=\bm{0}.$ (21)
Let $j*m==N$, we have:
$\lVert e_{m*j}\rVert_{2}$ $\leq\lVert e_{0}\rVert_{2}+\sum_{k=0}^{j-1}O(\max_{l\in\llbracket k*m,(k+1)*m-1\rrbracket}(t_{l+1}-t_{l})^{p+1})$ (22)
$=\sum_{k=0}^{j-1}O(\max_{l\in\llbracket k*m,(k+1)*m-1\rrbracket}(t_{l+1}-t_{l})^{p+1})$ (23)
$=\sum_{k=0}^{j-1}(\max_{l\in\llbracket k*m,(k+1)*m-1\rrbracket}(t_{l+1}-t_{l}))O(\max_{l\in\llbracket k*m,(k+1)*m-1\rrbracket}(t_{l+1}-t_{l})^{p})$ (24)
$\leq\sum_{k=1}^{j-1}(T-\epsilon)O(\max_{l\in\llbracket k*m,(k+1)*m-1\rrbracket}(t_{l+1}-t_{l})^{p})$ (25)
$\leq\sum_{k=1}^{j-1}(T-\epsilon)O((\Delta t)^{p})$ (26)
$=O((\Delta t)^{p})$ (27)
which completes the proof. Eq. 24 and Eq. 25 demonstrate that our method has a smaller error upper bound. ∎
Table 3: MSE Loss of Feature Extracted by DINOv2 During LGP and ACD Stages.
Stage MSE(DINOv2($x^{\text{Image}}_{in}$),DINOv2($x^{\text{Predict}}$)) MSE(DINOv2($f_{\theta}(x_{t_{n+m}},t_{n+m})$),DINOv2($x^{\text{Predict}}$))
LGP 0.21 0.26
ACD 0.0022 4.09e-5
Appendix B Discussion
B.1 Discussion on LGP and ACD
We demonstrate the convergence of training at different stages based on PCM [32]. Let the data distribution used in the LGP and ACD phases be denoted as $p_{0}$, and the forward conditional probability path is defined as $\alpha_{t}\mathbf{x}_{0}+\sigma_{t}\bm{\epsilon}$. The intermediate distribution is then defined as ${p}_{t}(\mathbf{x})=({p}_{0}(\frac{\mathbf{x}}{\alpha_{t}})\cdot\frac{1}{\alpha_{t}})\ast\mathcal{N}(0,\sigma_{t})$. Similarly, the data distribution used for pretraining the diffusion model is denoted as ${p}^{\text{pretrain}}_{0}(\mathbf{x})$, and the corresponding intermediate distribution during the forward process is ${p}^{\text{pretrain}}_{t}(\mathbf{x})=({p}^{\text{pretrain}}_{0}(\frac{\mathbf{x}}{\alpha_{t}})\cdot\frac{1}{\alpha_{t}})\ast\mathcal{N}(0,\sigma_{t})$. This is reasonable because current large diffusion models are typically trained with more resources on larger datasets compared to those used for consistency distillation. We denote $\mathcal{T}^{\bm{\phi}}_{t\to s}$, $\mathcal{T}^{\bm{\theta}}_{t\to s}$, and $\mathcal{T}^{\phi^{\prime}}_{t\to s}$ as the flow operators corresponding to the pre-trained diffusion model, the flow operators corresponding to our consistency model, and the PF-ODE of the data distribution used for consistency distillation, respectively.
We first discuss the convergence of $\mathcal{L}_{\text{ACD}}^{adv}$. We have $f_{\theta}(x_{t_{n+m}},t_{n+m})\equiv f_{\theta}(\hat{x}_{t_{n}}^{\phi},t_{n})$, where $x_{t_{n+m}}\in p_{n+m}$ and $x_{t_{n}}\in p_{n}$. Consequently, we obtain:
$\mathcal{T}^{\bm{\theta}}_{t_{n+m}\to\epsilon}\#\mathbb{P}_{t_{n+m}}\equiv\mathcal{T}^{\bm{\theta}}_{t_{n}\to\epsilon}\mathcal{T}^{\bm{\phi}}_{t_{n+m}\to t_{n}}\#\mathbb{P}_{t_{n+m}}\,.$ (28)
Therefore, if $Dis(f_{\theta}(x_{t_{n+m}},t_{n+m}),f_{\theta}(\hat{x}_{t_{n}}^{\phi},t_{n}))=0$, we have $\mathcal{L}_{\text{ACD}}^{adv}=0$.
We discuss the convergence of $\mathcal{L}_{\text{LGP}}^{adv}$. We have:
$p_{0}\equiv\mathcal{T}^{\bm{\phi}^{\prime}}_{t_{n+m}\to 0}\#p_{t_{n+m}}\,.$ (29)
Therefore, we have
$Dis\left(\mathcal{T}^{\bm{\theta}}_{t_{n+m}\to\epsilon}\#p_{t_{n+m}}\middle\|p_{0}\right)\,$ (30)
$=$ $Dis\left(\mathcal{T}^{\bm{\theta}}_{t_{n+m}\to\epsilon}\#p_{t_{n+m}}\middle\|\mathcal{T}^{\bm{\phi}^{\prime}}_{t_{n+m}\to 0}\#p_{t_{n+m}}\right)$ (31)
Because $f_{\theta}(x_{t_{n+m}},t_{n+m})\equiv f_{\theta}(\hat{x}_{t_{n}}^{\phi},t_{n})$, we have:
$Dis\left(\mathcal{T}^{\bm{\theta}}_{t_{n+m}\to\epsilon}\#p_{t_{n+m}}\middle\|\mathcal{T}^{\bm{\phi}^{\prime}}_{t_{n+m}\to 0}\#p_{t_{n+m}}\right)\,$ (32)
$=$ $Dis\left(\mathcal{T}^{\bm{\phi}}_{t_{n+m}\to\epsilon}\#p_{t_{n+m}}\middle\|\mathcal{T}^{\bm{\phi}^{\prime}}_{t_{n+m}\to 0}\#p_{t_{n+m}}\right)$ (33)
$=$ $Dis\left(p_{0}^{\text{pretrain}}\middle\|p_{0}\right)$ (34)
Because $p^{\text{pretrain}}_{0}\not=p_{0}$, we have $\mathcal{L}_{\text{LGP}}^{adv}>0$.
We consider the input condition $x_{\text{in}}^{\text{Image}}$ for the diffusion model, which involves replicating the image condition across multiple frames to align with the frame count of the original video. The output of our consistency model is $f_{\theta}(x_{t_{n+m}},t_{n+m})$. During the LGP phase, our prediction target is $x^{\text{Predict}}=x_{0}$. During the ACD phase, our prediction target is $x^{\text{Predict}}=f_{\theta^{-}}(\hat{x}^{\phi}_{t_{n}},t_{n})$.
We extract features from these data using DINOv2 and compute the MSE loss of these features. As shown in Table 3, during the LGP phase, the difference between $x_{\text{in}}^{\text{Image}}$ and $x^{\text{Predict}}$ is minimal, indicating that our consistency model tends to predict multiple static images. During the ACD phase, the difference between $f_{\theta}(x_{t_{n+m}},t_{n+m})$ and $x^{\text{Predict}}$ is minimal, indicating that our consistency model tends to predict data generated by the pre-trained model. Although random noise is added to $x_{\text{in}}^{\text{Image}}$ in actual training, this does not fundamentally solve the issue. However, fortunately, using LGP in the early stage of model training can accelerate the convergence of our distillation model. Figure 1 demonstrates the effectiveness of using LGP initially.
B.2 Contributions
Here, we re-emphasize the key components of OSV and summarize the contributions of our work.
The primary motivation of this research is to expedite the sampling process for high-resolution image-to-video generation by leveraging the consistency model training paradigm. Previous methods, including Animate-LCM and SF-V, sought to harness the potential of consistency models in this demanding scenario but failed to deliver satisfactory outcomes. We systematically examine and dissect the limitations of these approaches from three distinct perspectives. Crucially, these methods largely represent direct extensions of techniques originally devised to accelerate text-to-image sampling, and their straightforward adaptation to image-to-video sampling introduces significant challenges. To address these issues, we broaden the design space and propose comprehensive solutions to overcome these limitations.
The OSV framework is built upon the decomposition of the training process into two distinct stages, each utilizing a tailored distillation method to ensure efficient and effective model training. In the second stage, we introduce a multi-step solving method that capitalizes on the teacher model to execute multiple reverse ODE processes, thereby enhancing prediction accuracy. As illustrated in Figure 5, this multi-step solving method not only accelerates training but also significantly improves the performance of the consistency model.
Furthermore, inspired by the inherent properties of consistency models, we propose a novel higher-order solver, termed TTS, to replace the conventional CFG method. Experimental evaluations substantiate the efficacy of TTS, with results demonstrating state-of-the-art image-to-video generation performance. Remarkably, our approach achieves this using only 8 H800 GPUs (with merely 2 H800 GPUs required in the second stage), underscoring the efficiency and effectiveness of the proposed method.
B.3 Removing CFG
We introduce CFG into the distilled model: $\hat{\Phi}(\mathbf{x}_{t_{n+m}},t_{n+m},c;\phi)={\Phi}(\mathbf{x}_{t_{n+m}},t_{n+m},c_{zero};\phi)+w*({\Phi}(\mathbf{x}_{t_{n+m}},t_{n+m},c;\phi)-{\Phi}(\mathbf{x}_{t_{n+m}},t_{n+m},c_{zero};\phi)).$ This means the model already has CFG during inference, and using the same CFG scale again during inference leads to exposure issues in the generated videos. Table 2f also shows that a smaller CFG scale does not significantly improve the video quality. Removing CFG not only speeds up the model generation but also improves the overall quality of the generated videos.
Appendix C Additional Experimental Settings
$\lambda^{LGP}$ and $\lambda^{ACD}$ are set to 0.1. In the Huber Loss, we set $c=0.001$.
We train the model with videos of 14 frames, and the test videos also consist of 14 frames.
We use TTS only when the step equals 1.
Appendix D Upsampling Module
As shown in Figure 9, the upsampling module is displayed. First, we increase the number of channels of the latent space features, and then upsample the latent space features using the PixelShuffle operation. We set $r=4$.
Figure 9: Upsampling Module. We design the upsampling module inspired by sub-pixel convolution [25].